OpenStudy (anonymous):

2x/3-4/x=5/3 I need help with the process of finding the answer

OpenStudy (anonymous):

\[(2x*x - 4*3)/3x = 5/3\] \[(2x^2 - 12) = (5/3)*3x\] \[2x^2 - 12 = 5x\] \[2x^2 - 5x - 12 = 0\]